OpenStudy (anonymous):

A juggler is performing an act by juggling several balls. The juggler throws the balls up at an initial height of 3.9 feet with a speed of 14.8 feet per second. If the juggler did not catch a ball, about how long will it take the ball to hit the floor? 1.45 secs 1.14 secs 0.28 secs 0.21 secs

OpenStudy (anonymous):

Do..u knw kinematics equations of motion

OpenStudy (anonymous):

um.....no

OpenStudy (lgbasallote):

^^that sounds like physics

OpenStudy (anonymous):

V = u + at

OpenStudy (anonymous):

um... i thought it was H(t)= -16t^2+vt+s

OpenStudy (anonymous):

first we need to convert feet to meter

OpenStudy (anonymous):

ok...whats next

OpenStudy (anonymous):

then the time it takes to reach the maximum height

OpenStudy (anonymous):

wait why do i change it from feet to meters?

OpenStudy (anonymous):

because we are going to use a=g=-9.81m/s^2

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

whats the a and g? sorry

OpenStudy (anonymous):

that is accleration

OpenStudy (anonymous):

ok.......walk me through

OpenStudy (anonymous):

convert 3.9 feet into meter

OpenStudy (anonymous):

about one meter?

OpenStudy (anonymous):

14.8 feet per second in meter per second?

OpenStudy (anonymous):

oh okay 4 meters and something

OpenStudy (anonymous):

3.9 feet = 1.18872m and 14.8 feet per second=4.51104 meter per second

OpenStudy (anonymous):

now,v=u+at 0=4.51 + (-9.81)t t=0.46seconds

OpenStudy (anonymous):

Again, v^2=u^2+2as 0^2=(4.51)^2 + 2(-9.81)s s=1.04 m

OpenStudy (anonymous):

So, maximum height = 1.19+1.04 m = 2.23m

OpenStudy (anonymous):

Now, s=ut+1/2 at^2 2.23 = 0*t + 1/2 (9.81) t^2 t=0.67 seconds

OpenStudy (anonymous):

So, total time = 0.46+0.67 =1.13 seconds

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ohh okay.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome